Q(1) [10 Marks] [CLO2,C3] (a) The time required, in hours, to repair a car is a r.v. X with pdf fX(x) = c(4x − x 2 ), where 0 ≤ x ≤ 4. (i) Find the value of the constant c. (ii) Find the probability that for a car which arrives now at the garage, the amount of time needed to get repaired will be (a) at least one but less than three hours; (b) more than two hours.

(b) Let X be a r.v. whose pdf is: fX(x) = √ 1 2π exp ( − (x ^2) /2 ) , where −∞ < x < ∞. Determine the conditional pdf fX(x||x| ≤ 1). 


1
Expert's answer
2021-06-07T15:41:35-0400

i.

"\int_0^4c(4x-x^2)dx=1"

"c(2x^2-\frac{1}{3}x^3)|_0^4=1"

"c(32-\frac{64}{3})=1"

"32-21\frac{1}{3}=\frac{1}{c}"

"\frac{32}{3}=\frac{1}{c}"

"32c=3"

"c=\frac{3}{32}"

iia. "P(1\le x<3)"

"P(1\le x<3)=\int_1^3\frac{3}{32}(4x-x^2)dx"

"=\frac{3}{32}(2x^2-\frac{1}{3}x^3)|_1^3"

"=\frac{3}{32}[(18-\frac{27}{3})-(2-\frac{1}{3})]"

"=\frac{3}{32}(9-\frac{5}{3})"

"=\frac{3}{32}\times\frac{22}{3}"

"=\frac{11}{16}"

iib. "P(x>2)"

"P(x>2)=\int_2^4\frac{3}{32}(4x-x^2)dx"

"=\frac{3}{32}(2x^2-\frac{1}{3}x^3)|_2^4"

"=\frac{3}{32}[(32-\frac{64}{3})-(8-\frac{8}{3})]"

"=\frac{3}{32}(\frac{32}{3}-\frac{16}{3})"

"=\frac{3}{32}\times\frac{16}{3}"

"=\frac{1}{2}"

b. The pdf is not clear but it looks like a standard normal distribution

"fX(x||x| ≤ 1)=\frac{f(X(x)}{f(x)}" . |x|≤ 1

"-\infin<x<\infin\cap|x|\le 1=-1\le x\le 1"

"f(x)=\int_{-1}^1\frac{1}{\sqrt{2\pi}}e^{(-\frac{x^2}{2})} dx"

Which is equivalent to "P(-1\le Z\le 1)"

From Z tables or excel formula =NORM.S.DIST(1,TRUE)-NORM.S.DIST(-1,TRUE)),"P(Z\le1)-P(Z\le -1)=0.682689492"

Thus,

"fX(x||x| ≤ 1)=\frac{\frac{1}{\sqrt{2\pi}}e^{(-\frac{x^2}{2})}}{0.682689492}" ,"-1\le x \le 1"

Simplify by converting 0.682689492 as a fraction

"0.682689492=\frac{170672373}{250000000}"

"fX(x||x| ≤ 1)=\dfrac{1953125{\cdot}2^\frac{13}{2}\mathrm{e}^{-\frac{x^2}{2}}}{170672373\sqrt{{\pi}}},-1\le x \le 1"